Find the coordinates of the points of intersection of y= 2/x -5 and y= 4x +2
Answer by stanbon(75887)

Find the coordinates of the points of intersection of y= 2/x -5 and y= 4x +2
--------------
Solve::
2/x - 5 = 4x+2
----------------------
Multiply thru by "x" to get:
2 - 5x = 4x^2 + 2x
------
4x^2 + 7x - 2 = 0
-----
4x^2 +8x -x - 2 = 0
4x(x+2)-(x+2) = 0
(x+2)(4x-1) = 0
------
x = -2 or x = 1/4
----------------
Solve for "y" if y = 4x+2
If x = -2, y = -6
If x = (1/4), y = 3
Ans:: (-2,-6) and (1/4,3)
----------------
Cheers,
Stan H.
-----------